The erection of a tower takes 24 men to complete in 15 days. How many days will it take 9 men to erect if they work at the same rate?
Answer by ptaylor(2198)

If 24 men can complete the tower in 15 days, there are 24*15=360 man-days of labor involved in the project
In other words, 1 man can complete the project in 360 days.
9 men can complete the job in 360/9=40 days
And, of course, it takes 24 men 360/24=15 days to complete the job
Another way:
Let x=number of days it takes 9 men to do the job
The men work at the rate of 1/360 of the project per day
so, (9/360)*x=1 (1 tower, that is)
9x=360
x=40 days
Hope this helps---ptaylor